OpenStudy (luigi0210):

find the anti-derivative of \[f \prime \prime(x) = 2e ^{3}+3\sin t\], f(0)=0, f(pi)=0

OpenStudy (flywalker):

in my way: f'(x)= (2/3)*e^3-3cost+c and you are doing the same thing to go to f(x) and then (u find 1/f(x) and you find the derivative but i dont understand what is f(pi) to be sure if i am right! tell me to right the full explanation!

zepdrix (zepdrix):

@Luigi0210 do you understand how to find the anti-derivative of the \(2e^3\) portion?

zepdrix (zepdrix):

Unless you made a typo, there appears to be no \(t\) involved. So it is simply a constant. What is the anti-derivative of a constant?

OpenStudy (luigi0210):

I made a mistake, sorry. It's suppose to be 2e^t

OpenStudy (flywalker):

and i thing i made a big mistake

zepdrix (zepdrix):

oh ok :)

zepdrix (zepdrix):

Do you remember the derivative of \(\large e^x\)? It gives us the same thing back right? That process will be the same in reverse. So the anti-derivative of \(\large e^t\) is \(\large e^t\). That part is pretty straightforward yes? Just ignore the 2 in front, it has no effect on the anti-differentiation process.

zepdrix (zepdrix):

\[\large f''(x)=2e^t+3\sin t\] The anti-derivative of \(\sin t\) will be \(-\cos t\). If you're unclear about that, take the derivative of \(-\cos t\) and see what it gives you. You should end up with \(\sin t\). So we get something like this when we anti-differentiate once,\[\large f'(x)=2e^t-3\cos t+c\]

OpenStudy (luigi0210):

Yup, I got that far. But what I don't understand is finding the next anti-derivative

zepdrix (zepdrix):

Hmm it will be a similar process, but now we have a constant that showed up. Do you remember the anti-derivative of a constant? :)

zepdrix (zepdrix):

Or maybe an easier way to ask that would be, do you remember the anti-derivative of \(1\)?

OpenStudy (luigi0210):

it'll be 2e^t-3sint+Cx+D

zepdrix (zepdrix):

yes good c:

OpenStudy (luigi0210):

so how do I solve? Because now I have two variables..

OpenStudy (flywalker):

it's f(x) not f(t) you don't have f(x)=2e^t-3sint+Cx+D but f(x)=2(x**2/2)e^t-3sint+Cx+D actually 2e^t is a number!

zepdrix (zepdrix):

Oh I'm sure that was a typo :) lol

zepdrix (zepdrix):

Probably should be all t's or all x's.

OpenStudy (luigi0210):

yea, sorry about that

zepdrix (zepdrix):

Oh I was being sloppy earlier and writing f(x) also, lolol just noticed that XD

zepdrix (zepdrix):

So you were able to figure this out, \[\large f(t)=2e^t-3\sin t+Ct+D\] Now to solve for C and D, use your initial conditions,\[f(0)=0, \qquad \qquad f(\pi)=0\]

OpenStudy (flywalker):

if u put 0 for x and 0 for t you'll find find the one variable relative to the other!!

OpenStudy (flywalker):

pi= π???Pls tell

zepdrix (zepdrix):

\[\large f(0)=0 \qquad \rightarrow \qquad 0=2e^0-3 \sin 0+ C \cdot 0 +D\]See how that works Luigi? c:

OpenStudy (luigi0210):

yes, pi = \[\pi \]

zepdrix (zepdrix):

Hint,\[\large e^0=1\]

OpenStudy (luigi0210):

Oh right I forgot. So D=2

OpenStudy (flywalker):

ty!now you will change f(t)=y and you will have y=2et−3sint+Ct+D (keep in mind that you have to find C and D) after that have to so lve for t!!and that's anti f(t)!!after you find the second derivative of the new function

OpenStudy (flywalker):

but i think it's really hard to solve for t!!

OpenStudy (flywalker):

it is ok?

OpenStudy (flywalker):

(is it)*

OpenStudy (luigi0210):

all we need to do is just find the equation

OpenStudy (flywalker):

yes if you find the anti f(t) then it's easy!the problrm is that you have sin t and \[e^{t}\]